### v1

@@ 6 Discussion
The country-level results reveal a core-periphery pattern. Luxembourg (+60%), Belgium (+40%), and Netherlands (+39%) — small, open economies at the geographic heart of the eurozone — benefited most from euro adoption. These countries serve as logistics and financial hubs with trade-to-GDP ratios exceeding 100%, [^3] World Bank data show trade-to-GDP ratios of approximately 400% for Luxembourg, 160% for Belgium, and 150% for the Netherlands in the early 2000s. making them uniquely sensitive to transaction cost reductions. Peripheral economies show smaller effects: Greece (+11%), Portugal (+13%), and Finland (+15%). These countries have smaller domestic markets, fewer natural trading partners within the eurozone core, and greater geographic distance from the European economic center.
@@ 6.2 Dynamic Heterogeneity: Do Effects Evolve Over Time?
1999-2003 High-CATE 0.231 +25.9% [0.224, 0.237] 265
Low-CATE 0.182 +19.9% [0.179, 0.184] 260
All pairs 0.206 +22.9% [0.202, 0.210] 525
2004-2008 High-CATE 0.131 +14.0% [0.110, 0.153] 265
Low-CATE 0.220 +24.6% [0.201, 0.240] 260
All pairs 0.175 +19.2% [0.160, 0.190] 525
2009-2015 High-CATE 0.177 +19.4% [0.164, 0.191] 371
Low-CATE 0.142 +15.3% [0.134, 0.150] 364
All pairs 0.160 +17.3% [0.152, 0.168] 735
@@ 7 Counterfactual Analysis: Non-Eurozone EU Countries
Sweden +19.9 [+4.1, +38.1] 0.14 -14.9 +78.3
Denmark +19.4 [+4.0, +37.2] 0.18 -15.1 +74.3
United Kingdom +23.9 [+13.9, +34.8] 0.13 +7.5 +81.9
@@ 7.1 Counterfactual Support and Validity
The United Kingdom shows the largest predicted effect (+23.9%, std 0.13, range +7.5% to +81.9%), followed by Sweden (+19.9%, std 0.14, range $14.9\%$ to +78.3%) and Denmark (+19.4%, std 0.18, range $15.1\%$ to +74.3%). These differences reflect each country’s trade structure and existing integration with eurozone partners. Figure 12 summarizes the counterfactual trade trajectories for all three countries, showing actual trade with eurozone partners versus the predicted counterfactual under euro adoption.
@@ 7.2 Sweden
Sweden’s predicted average effect is +19.9% (std 0.14), with effects ranging from $14.9\%$ to +78.3% across partners. This masks substantial partner-level heterogeneity. Figure 13 shows the predicted effect by trading partner.
@@ 7.3 Denmark
Denmark shows a predicted effect of +19.4% (std 0.18), with effects ranging from $15.1\%$ to +74.3% across partners. Figure 16 shows the partner-level breakdown.
@@ 7.4 United Kingdom
The United Kingdom shows the largest predicted effect (+23.9%, std 0.13) with uniformly positive effects across all partners, ranging from +7.5% to +81.9%. Figure 19 shows the partner-level breakdown.
@@
Our UK estimate (+23.9%) is comparable to Saia (2017), who uses the synthetic control method and finds +16%. The difference likely reflects: (1) our method captures partner-level heterogeneity; and (2) SCM constructs a single synthetic counterfactual, while we estimate effects conditional on pair characteristics.
@@ 7.6 Illustrative Counterfactual Exercise
For the United Kingdom, the predicted +23.9% effect represents the largest potential gain among the three non-eurozone EU members. The uniformly positive effects across all partners (minimum +7.5%) suggest that UK-eurozone trade may have faced substantial currency friction that euro adoption could have reduced. Post-Brexit, this counterfactual becomes moot, but the analysis suggests that the UK’s decision to remain outside the eurozone may have carried some trade costs during its EU membership—though the magnitude is uncertain given the structural differences between the UK and eurozone economies.
For Sweden, the predicted +19.9% effect is substantial but comes with considerable partner-level variation (ranging from $14.9\%$ to +78.3%). Sweden’s largest predicted gains would come from smaller eurozone economies (Luxembourg, Austria, Belgium) rather than its major trading partners. This pattern suggests that Sweden’s existing trade relationships with Germany and Finland may already benefit from low currency friction, potentially limiting the marginal gains from formal euro adoption.
For Denmark, the predicted +19.4% effect is similar to Sweden’s, but the krone’s peg to the euro already captures much of the currency stability benefit. Denmark’s near-zero predicted effects with some peripheral partners suggest that for certain pairs, the benefits of reduced transaction costs may be modest. Denmark’s current arrangement—maintaining the peg without full adoption—may represent a reasonable middle ground, though this is speculative.
@@ 8 Conclusion
The results suggest a positive relationship between euro adoption and bilateral trade, but the magnitude varies substantially across country pairs. Our preferred average effect estimate is around 15% (14.7% after fixed effects correction), consistent with gravity benchmarks; the 24% naive estimate reflects heterogeneity-weighted averages that give more weight to high-effect pairs. This heterogeneity explains why prior studies using different samples and methods have produced such divergent estimates. The variation is not methodological noise; it appears to reflect genuine differences in how the euro affected different trading relationships.
Core eurozone pairs with strong pre-existing trade relationships experienced larger effects, while peripheral pairs saw smaller gains. The euro amplified existing trade relationships rather than creating new ones, with pre-euro trade intensity and GDP as key drivers of heterogeneity. Counterfactual analysis suggests non-eurozone EU members would have experienced moderate effects: UK (+23.9%), Sweden (+19.9%), Denmark (+19.4%).
@@
Implications for Prospective Euro Members
Our findings have direct relevance for the seven EU members that have not yet adopted the euro: Bulgaria, Czech Republic, Denmark, Hungary, Poland, Romania, and Sweden. Each faces a distinct cost-benefit calculus shaped by their current trade patterns, exchange rate arrangements, and economic structure.
@@
Central European economies (Poland, Czech Republic, Hungary) have deep trade integration with Germany and other core eurozone members, suggesting substantial potential gains from adoption. However, these countries also maintain flexible exchange rates that have served as shock absorbers during economic downturns. The trade-off between transaction cost reduction and monetary policy flexibility is particularly acute for these economies.
Southeastern European economies (Bulgaria, Romania) have weaker existing trade ties with the eurozone core, suggesting more modest potential gains. Bulgaria’s currency board arrangement already provides exchange rate stability with the euro, capturing some benefits without full adoption. Romania’s floating exchange rate implies larger potential gains but also greater adjustment costs.
Nordic holdouts (Denmark, Sweden) present interesting cases. Denmark’s euro peg already captures most currency stability benefits, and our counterfactual analysis suggests modest marginal gains from formal adoption. Sweden’s managed float provides more flexibility, and our estimates suggest larger potential gains—but Sweden has twice rejected euro adoption in referenda, reflecting preferences for monetary policy autonomy that our trade-focused analysis cannot capture.
Looking Forward
The eurozone of 2025 differs substantially from the eurozone of 1999. The institutional framework has evolved through the European Stability Mechanism, banking union, and enhanced fiscal coordination. The membership has expanded from 11 to 20 countries. Global trade patterns have shifted with the rise of China and the fragmentation of supply chains following COVID-19 and geopolitical tensions.
These changes suggest caution in extrapolating our historical estimates to future adoption decisions. Countries considering adoption today face a more mature, institutionally developed eurozone—but also one that has weathered severe crises and faces ongoing challenges in coordinating fiscal and monetary policy across diverse economies. The trade benefits we document are real, but they represent only one input into a complex policy decision that involves sovereignty, macroeconomic stability, and political economy considerations beyond the scope of this analysis.
@@
What our analysis does establish is that the euro’s trade effects are heterogeneous, predictable, and substantial for the right country pairs. Countries with strong existing trade ties to the eurozone core, floating exchange rates that create genuine currency friction, and economic structures complementary to European supply chains stand to gain most. Those with weak eurozone trade ties, existing currency stability arrangements, or trade oriented toward non-eurozone partners may find the trade case for adoption less compelling. The causal forest framework provides a tool for making these assessments with partner-level granularity, moving beyond the single-number estimates that have dominated policy debates.
Limitations and Future Research
Our analysis focuses on aggregate bilateral trade flows, but the euro’s effects may vary across product categories. High-technology goods with complex supply chains and frequent cross-border transactions might benefit more from reduced currency friction than homogeneous commodities traded on global markets. Similarly, differentiated products where invoicing currency matters could show larger effects than standardized goods with established pricing conventions. Future research using product-level bilateral trade data (e.g., SITC or HS classifications from sources like CEPII BACI or UN Comtrade) could decompose the aggregate effects we document into product-specific components, testing whether the euro disproportionately boosted trade in technology-intensive or differentiated goods. Such analysis would shed light on the mechanisms through which currency unions affect trade and provide more granular guidance for countries with specialized export structures.
@@ A.1.4 Alternative Outcome Measures
To assess whether our results are sensitive to the choice of outcome variable, we re-estimate the causal forest using alternative measures of bilateral trade. Table A.2 shows results for log exports only and log imports only. The baseline effect on log bilateral trade is 11.9%. Effects are similar for exports (+15.3%) and imports (+10.1%), with overlapping confidence intervals. This symmetry is reassuring: exports in our data measure the reporter’s outbound trade while imports measure inbound trade, so similar effects suggest the euro boosted trade flows in both directions rather than favoring one side. The slightly larger point estimate for exports is not statistically distinguishable from the imports effect.
@@
Outcome Measure ATE Effect 95% CI N
Log Bilateral Trade (Baseline) 0.183 +20.1% [0.068, 0.299] 1,951
Log Exports Only 0.260 +29.7% [0.066, 0.454] 1,951
Log Imports Only 0.087 +9.1% [-0.091, 0.265] 1,951
Trade Growth (YoY) 0.008 +0.8 pp [-0.026, 0.041] 1,860
Log Trade Intensity 0.241 +27.3% [0.087, 0.395] 1,951
Notes: All models estimated using CausalForestDML on EU15 bilateral trade data, 1988-2015. Effect shows percentage change for log outcomes and percentage points for growth outcomes. Trade growth is year-over-year change in log bilateral trade. Trade intensity is bilateral trade divided by GDP product.
@@ A.1.5 Sensitivity to Unobserved Confounding
Following Oster (2019), we assess how much selection on unobservables would be required to explain away our results. Table A.3 presents the analysis. Panel A shows coefficient stability across specifications: the coefficient increases from near-zero without controls to 0.38 with GDP and year fixed effects, then falls to 0.15 with two-way fixed effects. Panel B shows bias-adjusted estimates for different assumptions about the degree of selection on unobservables ($\delta$) and the maximum R-squared ($R_{max}$). The analysis suggests that unobserved confounders would need to be substantially more important than observed confounders to reduce the effect to zero. Since $\delta<0$ (the coefficient increases with controls), omitted variables appear to bias the effect downward rather than upward.
@@
No controls -0.004 -0.4% 0.000 2,189
GDP controls 0.015 +1.6% 0.740 2,189
GDP + Year FE 0.380 +46.3% 0.790 2,189
Two-way FE 0.151 +16.3% 0.627 2,189
@@
0.5 0.383 0.407 0.431
1.0 0.385 0.434 0.483
1.5 0.388 0.461 0.534
2.0 0.390 0.488 0.585
Notes: Panel A shows coefficient stability across specifications. Panel B shows bias-adjusted coefficients ($\beta^{*}$) following Oster (2019). $\delta$ is the ratio of selection on unobservables to observables. $R_{max}$ is the hypothetical R-squared if all relevant variables were included. For $R_{max}=1.0$, $\delta=-0.26$ would be needed to explain away the effect. Since $\delta<0$, the coefficient increases with controls, suggesting omitted variables bias the effect downward.
@@ A.1.6 Time Stability
Table A.4 shows how the EU15 estimate evolves as we extend the sample year by year from 2007 to 2019. The estimates remain remarkably stable, ranging from 20.5% to 24.5% across all time windows. This stability suggests our main findings are not sensitive to the choice of end year and are robust to the inclusion of crisis and post-crisis periods.
@@ A.1.7 Extending to EU28: The Puzzle
A natural question is whether our results extend to the full EU28, including countries that joined the EU after 2004 and adopted the euro during the 2008–2015 period. Table A.5 shows how the estimated euro effect changes as we progressively add countries to the EU15 baseline. The results reveal a striking pattern: the ATE remains stable around 23–26% as we add Slovenia (2007), Cyprus (2008), and Malta (2008), but begins to decline with Slovakia (2009) and drops sharply with the Baltic states (Estonia 2011, Latvia 2014, Lithuania 2015). Adding Lithuania produces a negative estimate of $5.5\%$. By the time we include all EU28 members, the naive causal forest estimate falls to just 5.8% — a quarter of the EU15 estimate of 23.6%.
@@
EU15 2,609 1,364 1,245 0.212 23.6
+Slovenia (2007) 2,968 1,520 1,448 0.233 26.3
+Cyprus (2008) 3,312 1,676 1,636 0.218 24.4
+Malta (2008) 3,705 1,844 1,861 0.221 24.8
+Slovakia (2009) 4,127 2,009 2,118 0.208 23.1
+Estonia (2011) 4,538 2,153 2,385 0.194 21.4
+Latvia (2014) 4,994 2,255 2,739 0.137 14.7
+Lithuania (2015) 5,471 2,345 3,126 -0.057 -5.5
+Croatia (2023) 5,941 2,345 3,596 0.171 18.6
+Poland (never) 6,464 2,345 4,119 0.108 11.4
+Czech Republic (never) 6,976 2,345 4,631 0.004 0.4
+Hungary (never) 7,529 2,345 5,184 -0.092 -8.8
+Romania (never) 8,119 2,345 5,774 -0.014 -1.4
+Bulgaria (never) 8,694 2,345 6,349 0.047 4.8
Notes: Each row adds one country to the sample. Year in parentheses indicates euro adoption date. Data covers 1988–2019. ATE is the average treatment effect in log points.

### v2

@@ 6 Discussion
The country-level results reveal a core-periphery pattern. Luxembourg (+69%), Germany (+50%), Belgium (+47%), and Netherlands (+47%) — small, open economies at the geographic heart of the eurozone — benefited most from euro adoption. These countries serve as logistics and financial hubs with trade-to-GDP ratios exceeding 100%, [^3] World Bank data show trade-to-GDP ratios of approximately 400% for Luxembourg, 160% for Belgium, and 150% for the Netherlands in the early 2000s. making them uniquely sensitive to transaction cost reductions. Peripheral economies show smaller effects: Greece (+10%), Portugal (+16%), and Finland (+15%). These countries have smaller domestic markets, fewer natural trading partners within the eurozone core, and greater geographic distance from the European economic center.
@@ 6.2 Dynamic Heterogeneity: Do Effects Evolve Over Time?
1999-2003 High-CATE 0.242 +27.4% [0.235, 0.248] 265
Low-CATE 0.218 +24.4% [0.215, 0.221] 260
All pairs 0.230 +25.9% [0.226, 0.234] 525
2004-2008 High-CATE 0.154 +16.6% [0.125, 0.183] 265
Low-CATE 0.230 +25.9% [0.216, 0.245] 260
All pairs 0.192 +21.1% [0.175, 0.208] 525
2009-2015 High-CATE 0.160 +17.3% [0.145, 0.175] 371
Low-CATE 0.189 +20.8% [0.180, 0.198] 364
All pairs 0.174 +19.0% [0.165, 0.183] 735
@@ 7 Counterfactual Analysis: Non-Eurozone EU Countries
Sweden +21.9 [+7.7, +38.1] 0.16 -14.9 +85.9
Denmark +19.1 [+4.8, +35.5] 0.19 -16.1 +85.7
United Kingdom +32.6 [+21.4, +44.9] 0.16 +10.3 +80.1
@@ 7.1 Counterfactual Support and Validity
The United Kingdom shows the largest predicted effect (+32.6%, std 0.16, range +10.3% to +80.1%), followed by Sweden (+21.9%, std 0.16, range $14.9\%$ to +85.9%) and Denmark (+19.1%, std 0.19, range $16.1\%$ to +85.7%). These differences reflect each country’s trade structure and existing integration with eurozone partners. Figure 12 summarizes the counterfactual trade trajectories for all three countries, showing actual trade with eurozone partners versus the predicted counterfactual under euro adoption.
@@ 7.2 Sweden
Sweden’s predicted average effect is +21.9% (std 0.16), with effects ranging from $14.9\%$ to +85.9% across partners. This masks substantial partner-level heterogeneity. Figure 13 shows the predicted effect by trading partner.
@@ 7.3 Denmark
Denmark shows a predicted effect of +19.1% (std 0.19), with effects ranging from $16.1\%$ to +85.7% across partners. Figure 16 shows the partner-level breakdown.
@@ 7.4 United Kingdom
The United Kingdom shows the largest predicted effect (+32.6%, std 0.16) with uniformly positive effects across all partners, ranging from +10.3% to +80.1%. Figure 19 shows the partner-level breakdown.
@@
Our UK estimate (+32.6%) is comparable to Saia (2017), who uses the synthetic control method and finds +16%. The difference likely reflects: (1) our method captures partner-level heterogeneity; and (2) SCM constructs a single synthetic counterfactual, while we estimate effects conditional on pair characteristics.
@@ 7.6 Illustrative Counterfactual Exercise
For the United Kingdom, the predicted +32.6% effect represents the largest potential gain among the three non-eurozone EU members. The uniformly positive effects across all partners (minimum +10.3%) suggest that UK-eurozone trade may have faced substantial currency friction that euro adoption could have reduced. Post-Brexit, this counterfactual becomes moot, but the analysis suggests that the UK’s decision to remain outside the eurozone may have carried some trade costs during its EU membership—though the magnitude is uncertain given the structural differences between the UK and eurozone economies.
For Sweden, the predicted +21.9% effect is substantial but comes with considerable partner-level variation (ranging from $14.9\%$ to +85.9%). Sweden’s largest predicted gains would come from smaller eurozone economies (Luxembourg, Austria, Belgium) rather than its major trading partners. This pattern suggests that Sweden’s existing trade relationships with Germany and Finland may already benefit from low currency friction, potentially limiting the marginal gains from formal euro adoption.
For Denmark, the predicted +19.1% effect is similar to Sweden’s, but the krone’s peg to the euro already captures much of the currency stability benefit. Denmark’s near-zero predicted effects with some peripheral partners suggest that for certain pairs, the benefits of reduced transaction costs may be modest. Denmark’s current arrangement—maintaining the peg without full adoption—may represent a reasonable middle ground, though this is speculative.
@@ 8 Conclusion
The results suggest a positive relationship between euro adoption and bilateral trade, but the magnitude varies substantially across country pairs. Our preferred average effect estimate is around 15% (14.1% after fixed effects correction), consistent with gravity benchmarks; the 29% naive estimate reflects heterogeneity-weighted averages that give more weight to high-effect pairs. This heterogeneity explains why prior studies using different samples and methods have produced such divergent estimates. The variation is not methodological noise; it appears to reflect genuine differences in how the euro affected different trading relationships.
Core eurozone pairs with strong pre-existing trade relationships experienced larger effects, while peripheral pairs saw smaller gains. The euro amplified existing trade relationships rather than creating new ones, with pre-euro trade intensity and GDP as key drivers of heterogeneity. Counterfactual analysis suggests non-eurozone EU members would have experienced moderate effects: UK (+32.6%), Sweden (+21.9%), Denmark (+19.1%).
@@
Our findings have direct relevance for the seven EU members that have not yet adopted the euro: Bulgaria, Czech Republic, Denmark, Hungary, Poland, Romania, and Sweden. Each faces a distinct cost-benefit calculus shaped by their current trade patterns, exchange rate arrangements, and economic structure. Central European economies (Poland, Czech Republic, Hungary) have deep trade integration with Germany and other core eurozone members, suggesting substantial potential gains from adoption, though these countries also maintain flexible exchange rates that have served as shock absorbers during economic downturns. Southeastern European economies (Bulgaria, Romania) have weaker existing trade ties with the eurozone core, suggesting more modest potential gains. Nordic holdouts (Denmark, Sweden) present interesting cases: Denmark’s euro peg already captures most currency stability benefits, while Sweden’s managed float provides more flexibility but our estimates suggest larger potential gains.
@@
The eurozone of 2025 differs substantially from the eurozone of 1999, with an evolved institutional framework, expanded membership from 11 to 20 countries, and shifted global trade patterns. These changes suggest caution in extrapolating our historical estimates to future adoption decisions. What our analysis does establish is that the euro’s trade effects are heterogeneous, predictable, and substantial for the right country pairs. Countries with strong existing trade ties to the eurozone core, floating exchange rates that create genuine currency friction, and economic structures complementary to European supply chains stand to gain most. The causal forest framework provides a tool for making these assessments with partner-level granularity, moving beyond the single-number estimates that have dominated policy debates.
Our analysis focuses on aggregate bilateral trade flows, but the euro’s effects may vary across product categories. Future research using product-level bilateral trade data could decompose the aggregate effects we document into product-specific components, testing whether the euro disproportionately boosted trade in technology-intensive or differentiated goods.
@@ A.1.4 Alternative Outcome Measures
To assess whether our results are sensitive to the choice of outcome variable, we re-estimate the causal forest using alternative measures of bilateral trade. Table A.2 shows results for five outcome specifications. The baseline effect on log bilateral trade is +24.0%. Effects are similar for exports (+17.3%) and imports (+16.7%), with overlapping confidence intervals. This symmetry is reassuring: exports in our data measure the reporter’s outbound trade while imports measure inbound trade, so similar effects suggest the euro boosted trade flows in both directions rather than favoring one side. The slightly larger point estimate for exports is not statistically distinguishable from the imports effect. Year-over-year trade growth shows a small effect (+1.1%), consistent with the level effects accumulating over time. Log trade intensity (trade normalized by GDP) shows a larger effect (+31.1%), suggesting the euro increased trade relative to economic size.
@@
Outcome ATE Effect 95% CI N
Log Bilateral Trade (Baseline) 0.215 +24.0% [0.091, 0.339] 1,911
Log Exports Only 0.160 +17.3% [-0.007, 0.326] 3,822
Log Imports Only 0.155 +16.7% [-0.018, 0.327] 3,822
Trade Growth (YoY) 0.011 +1.1 pp [-0.020, 0.043] 1,820
Log Trade Intensity 0.270 +31.1% [0.117, 0.424] 1,911
[…]
@@ A.1.5 Sensitivity to Unobserved Confounding
Following Oster (2019), we assess how much selection on unobservables would be required to explain away our results. Table A.3 presents the analysis. Panel A shows coefficient stability across specifications: the coefficient increases from near-zero without controls to 0.38 with GDP and year fixed effects, then falls to 0.16 with two-way fixed effects. Panel B shows bias-adjusted estimates for different assumptions about the degree of selection on unobservables ($\delta$) and the maximum R-squared ($R_{max}$). The analysis suggests that unobserved confounders would need to be substantially more important than observed confounders to reduce the effect to zero. Since $\delta<0$ (the coefficient increases with controls), omitted variables appear to bias the effect downward rather than upward.
@@
No controls 0.019 +1.9% 0.000 2,149
GDP controls 0.034 +3.5% 0.741 2,149
GDP + Year FE 0.382 +46.5% 0.789 2,149
Two-way FE 0.157 +17.0% 0.603 2,149
@@
0.5 0.385 0.408 0.431
1.0 0.387 0.433 0.479
1.5 0.390 0.459 0.528
2.0 0.392 0.484 0.576
Notes: Panel A shows coefficient stability across specifications. Panel B shows bias-adjusted coefficients ($\beta^{*}$) following Oster (2019). $\delta$ is the ratio of selection on unobservables to observables. $R_{max}$ is the hypothetical R-squared if all relevant variables were included. For $R_{max}=1.0$, $\delta=-0.28$ would be needed to explain away the effect. Since $\delta<0$, the coefficient increases with controls, suggesting omitted variables bias the effect downward.
@@ A.1.6 Time Stability
Table A.4 shows how the EU15 estimate evolves as we extend the sample year by year from 2007 to 2019. The estimates remain remarkably stable, ranging from 22.2% to 28.6% across all time windows. This stability suggests our main findings are not sensitive to the choice of end year and are robust to the inclusion of crisis and post-crisis periods.
@@ A.1.7 Extending to EU28: The Puzzle
A natural question is whether our results extend to the full EU28, including countries that joined the EU after 2004 and adopted the euro during the 2008–2015 period. Table A.5 shows how the estimated euro effect changes as we progressively add countries to the EU15 baseline. The results reveal a striking pattern: the ATE remains stable around 31.3–34.1% as we add Slovenia (2007), Cyprus (2008), and Malta (2008), but begins to decline with Slovakia (2009) and drops sharply with the Baltic states (Estonia 2011, Latvia 2014, Lithuania 2015). Adding Lithuania produces an estimate of just 0.9%. By the time we include all EU28 members, the naive causal forest estimate falls to just 4.3% — compared to the EU15 estimate of 28.6%.
@@
EU15 2,149 1,100 1,049 0.252 28.6
+Slovenia (2007) 2,928 1,520 1,408 0.272 31.3
+Cyprus (2008) 3,272 1,676 1,596 0.293 34.1
+Malta (2008) 3,665 1,844 1,821 0.269 30.9
+Slovakia (2009) 4,087 2,009 2,078 0.256 29.2
+Estonia (2011) 4,498 2,153 2,345 0.181 19.9
+Latvia (2014) 4,954 2,255 2,699 0.165 17.9
+Lithuania (2015) 5,431 2,345 3,086 0.009 0.9
+Croatia (2023) 5,901 2,345 3,556 0.192 21.2
+Poland (never) 6,424 2,345 4,079 0.093 9.7
+Czech Republic (never) 6,936 2,345 4,591 0.003 0.3
+Hungary (never) 7,489 2,345 5,144 -0.070 -6.7
+Romania (never) 8,079 2,345 5,734 0.031 3.2
+Bulgaria (never) 8,654 2,345 6,309 0.042 4.3
Notes: Each row adds one country to the sample. Year in parentheses indicates euro adoption date. Data covers 1995–2019. ATE is the average treatment effect in log points.